Can you please help solve this problem...instructions...complete the ordered paris for the equation y = 3x+3...( ,0), (1/3, ), (0, ), (-1/3, )
Answer by stanbon(75887)

complete the ordered paris for the equation y = 3x+3...( ,0), (1/3, ), (0, ), (-1/3, )
---------------
If y=0, 3x+3=0; x+1=0, x=-1 giving you (-1,0)
If x=1/3, y=3(1/3)+3, y=1+3, y=4 giving you (1/3,4)
If x=0, y=3*0+3; y=3, giving you (0,3)
If x=-1/3, y=3(-1/3)+3; y=-1+3, y=2, giving you (-1/3,2)
---------
Cheers,
Stan H.